### v2

@@ Abstract
Deep learning, a branch of artificial intelligence, is a computational model that uses multiple layers of interconnected units (neurons) to learn intricate patterns and representations directly from raw input data. Empowered by this learning capability, it has become a powerful tool for solving complex problems and is the core driver of many groundbreaking technologies and innovations. Building a deep learning model is a challenging task due to the algorithm’s complexity and the dynamic nature of real-world problems. Several studies have reviewed deep learning concepts and applications. However, the studies mostly focused on the types of deep learning models and convolutional neural network architectures, offering limited coverage of the state-of-the-art of deep learning models and their applications in solving complex problems across different domains. Therefore, motivated by the limitations, this study aims to comprehensively review the state-of-the-art deep learning models in computer vision, natural language processing, time series analysis and pervasive computing. We highlight the key features of the models and their effectiveness in solving the problems within each domain. Furthermore, this study presents the fundamentals of deep learning, various deep learning model types and prominent convolutional neural network architectures. Finally, challenges and future directions in deep learning research are discussed to offer a broader perspective for future researchers.
@@ 1 Introduction
Deep learning has revolutionized many applications across a variety of industries and research. The application of deep learning can be found in healthcare Shamshirband et al. (2021), smart manufacturing Wang et al. (2018b), robotics Pierson and Gashler (2017), cybersecurity Dixit and Silakari (2021) etc., solving challenging and complex problems such as disease diagnosis, anomaly detection, object detection and malware attack detection. Deep learning is a subset of machine learning that learns from data using artificial neural networks. An artificial neural network is a computational model that imitates the working principles of a human brain. The computational models are composed of an input layer which receives the input data, multiple processing layers that learn the representation of data and the output layer which produces the output of the model.
@@
Several studies have been conducted to discuss the concept and application of deep learning in the last few years, as listed in Table 1. The studies addressed or focused on several aspects of deep learning, such as types of deep learning models, learning approaches and strategies, convolutional neural network (CNN) architectures, deep learning applications and challenges. In Dong et al. (2021), the authors provided fundamentals of deep learning and highlighted different types of deep learning models, such as convolutional neural networks, autoencoder and generative adversarial networks. Then, the applications of deep learning in various domains are discussed, and some challenges associated with deep learning applications are presented. Another survey Talaei Khoei et al. (2023) provided a comprehensive analysis of supervised, unsupervised and reinforcement learning approaches and compared the different learning strategies such as online, federated and transfer learning. Finally, the current challenges of deep learning and future direction are discussed.
In Alzubaidi et al. (2021), the authors provided a comprehensive review of the popular CNN architectures used in computer vision tasks, highlighting their key features and advantages. Then, the applications of deep learning in medical imaging and the challenges are discussed. A similar survey is reported in Alom et al. (2019), where the different supervised and unsupervised deep learning models are highlighted, and the popular CNN architectures are compared and discussed. In another survey Pouyanfar et al. (2018), the authors focused on the applications of deep learning in computer vision, natural language processing and speech and audio processing. The different types of deep learning models are also discussed. In Sarker (2021), the authors focused on the different types of deep learning models and provided a summary of deep learning applications in various domains.
@@
Dong et al. (2021) A short review of the fundamentals of DL networks and discusses different types of neural networks, DL applications and challenges. Lack of analysis of CNN architectures and limited coverage of deep learning fundamentals.
Talaei Khoei et al. (2023) Discusses the learning approaches (supervised, unsupervised and reinforcement learnings), learning strategies, and DL challenges Lack of fundamentals of deep learning, CNN architectures and DL applications.
Alzubaidi et al. (2021) Discusses different types of DL networks, CNN fundamentals and architectures, DL challenges and medical imaging applications Limited discussion on DL applications such as natural language processing and time series analysis.
Alom et al. (2019) A short review of the fundamentals of neural networks and discusses different types of DL networks, CNN architectures and applications. Limited discussion on DL applications and no discussion of DL challenges.
Pouyanfar et al. (2018) Discusses different types of DL networks and DL applications and challenges Lack of analysis of CNN architectures and limited coverage of deep learning fundamentals.
Sarker (2021) Discusses different types of DL networks and provides a summary of DL applications Lack of fundamentals of deep learning, analysis of CNN architectures and limited discussion on DL applications.
@@ 2.1 Layers
A deep learning model is characterized by having numerous hidden layers. The hidden layers are responsible for learning and extracting complex features from the input data. A hidden layer is composed of an arbitrary number of neurons which serves as the fundamental building block of a neural network as shown in Fig. 1. A neuron consists of an arbitrary number of inputs, each associated with a weight, which controls the flow of information into the neuron during the forward pass. The flow of information, or forward pass, involves the computation of summation of the weighted input, followed by the application of a transformation function to the weighted sum. Consider a neuron $z_{i}$ at layer $l$, receives an input vector $a_{j}^{(}l-1)$, the computation of the neuron is defined as
$z_{i}^{l}=\sum_{j=0}^{d}w_{i,j}^{l-1}\cdot a_{j}^{l-1}$
$a_{i}^{l}=g(z_{i}^{l})$
where $w_{ij}$ is the set of weights connecting the inputs to the neuron and g is the transformation function also known as activation function. A hidden layer wherein each neuron is connected to all neurons of the previous layer is known as fully-connected layer.
@@
The key aspect of deep learning lies in its ability to automatically extract hierarchical features from the input data. This automatic feature extraction is performed by two specialized layers called convolutional and pooling layers. In a convolutional layer, each neuron is connected only to a local region of the input data, and the weights are shared across the input data. The weight-sharing not only significantly reduces the number of parameters of the neural network, but also allows the network to learn the same features across different spatial locations in the input LeCun et al. (2015). Fig. 2 illustrates a convolutional layer applies $3\times 3$ filter on two-dimensional image consisting of $9\times 9$ pixels. The layer convolves the input data by moving the filter across the whole input pixels, producing a set of output values called feature map. The computation of a convolutional layer $l$ is defined as
$z_{i,j,d}^{l}=\sum_{m=0}^{k_{1}}\sum_{n=0}^{k_{2}}w_{m,n,d}^{l-1}\cdot a_{i+m,j+n,d}^{l-1}$
$a_{i,j,d}^{l}=g(z_{i,j,d}^{l})$
where $w_{m,n,d}$ is the weight of $k_{1}\times k_{2}$ filter and $a_{i,j,d}^{l-1}$ is the input pixel.
@@
Pooling layers are commonly applied after successive convolutional layers to progressively reduce the spatial dimensions of the feature maps. The spatial reduction is performed by computing the summary of the local regions in the feature maps. Two common pooling operations are computing the maximum and the average of the local regions. This reduces the number of parameters while providing translation-invariant features LeCun et al. (2015).
@@ 2.2 Attention Mechanisms
where $g$ is a composite function that performs a sequence of operations to generate the attention or the weights and f applies the generated attention $g(x)$ on the input $x$, $f=g(x)x$.
@@
For instance, the squeeze-and-excitation (SE) attention generates the attention through five consecutive operations Hu et al. (2019). First, the input is vectorized using global average pooling. Then the vector is passed to two fully-connected layers, where the first one with ReLU activation and the second one with sigmoid activation. SE attention was a pioneer in channel attention. The attention module assigns varying weights to the channels of the feature maps. SE attention suffers from computational cost and the use of global average which may cause information loss at the spatial level. Several efforts have been made to improve SE attention. GSoP attention performs $1\times 1$ convolution on the feature maps to reduce the number of channels, and then computes the pairwise channel correlation which is used to generate the weights Gao et al. (2018). ECA attention replaced the fully-connected layers with 1d convolution to reduce the number of parameters and the computational cost Wang et al. (2020a).
@@
Temporal attention is an attention module that focuses on specific time steps in a sequence of data such time series and video (sequence of images). In video processing such as recognizing human actions, temporal attention is used to focus on key frames at different point in time that contains crucial information for predicting the ongoing activity. Temporal adaptive module (TAM) is a temporal attention that can focus on short-term (local) information and global context information of the data Liu et al. (2021b). The composite function consists of local branch for generating attention weights and global branch for generating channel-wise adaptive kernel. First, the input feature map is squeezed using global average pooling to reduce the computational cost. Subsequently, the local branch executes two 1D convolution operations, with the first convolution using ReLU, and sigmoid activation for the second to generate the local weights. The local weights are then multiplied with the featuremap. Meanwhile, the global branch is composed of two fully-connected layers, with the first layer using ReLU and second layer employing softmax function to generate the adaptive kernel (weights). Self-attention is a form of temporal attention, initially proposed for machine translation to enable the deep learning models to attend different words in a sequence relative to other words Vaswani et al. (2023). The attention module has become a fundamental building block in various natural language processing applications. To generate the attention weights, the input (word embeddings) is transformed by linear projection to compute query, key and value. Then, dot product between query and key is computed, and the resultant is normalized by the square root of the size of the key. Finally, the attention weights are obtained by applying the softmax function. Self-attention is the fundamental building block of the Transformer architecture, a key deep learning model in natural language processing.
@@
Spatial attention focuses on specific regions or spatial location of the input data, enabling the deep learning models to selectively emphasize and ignore certain features. In the context of computer vision, spatial attention is crucial in capturing the spatial relationships and context within an image for accurate prediction. Attention gate is a spatial attention that can identify and focus the salient regions and suppress feature responses of the insignificant ones. The composite function consists of ReLU activation followed by $1\times 1$ convolution to reduce channel dimension of the feature maps to a singular feature map. Finally, sigmoid is applied to the feature map to generate attention weights Oktay et al. (2018). The self-attention in the standard Transformer is not effective in handling image data due to its inherent sequential processing nature and lacks the ability to capture spatial dependencies and local patterns. To address this limitation, the Vision Transformer (ViT) treats images as a sequence of non-overlapping patches. A similar computational pipeline is used to generate the attention weights, the sequence of patches is transformed by linear projection to compute the query, key and value Dosovitskiy et al. (2021). The same operations are employed to generate the attention weights. Self-attention is computationally costly due to its quadratic complexity especially dealing with image data. To reduce the complexity, two learnable linear layers, independent of the input data are adopted as the key and value vectors Guo et al. (2023).
@@ 2.3 Activation Functions
The role of the activation function is to transform the weighted sum into a more classifiable form. This is crucial to the learning behaviour of the deep learning model, generating non-linear relationships between the input and the output of the model. The activation function, combined with many hidden layers allow the neural network to approximate highly complex, non-linear functions. Many activation functions are available for use in neural networks, and some of the functions are shown in Fig. 3 - Fig 5. The figures show the plot of the three popular activation functions. The sigmoid is a classic example of activation function which is used in logistic regression. Sigmoid activation function maps the weighted sum to a value in the range of 0 and 1 which can be used for classification. Hyperbolic tangent is another popular choice of bounded activation function which produces an output between -1 and 1. Hyperbolic tangent has a stronger gradient, hence the neural network training often converges faster than sigmoid LeCun et al. (2012). For many years, sigmoid and hyperbolic tangent are the commonly used activation functions. Nevertheless, the activation functions suffer from vanishing gradient problem, hindering the efficient training of deep neural networks with many layers Hochreiter et al. (2001).
It was shown that neural networks with unbounded activation functions have the universal approximation property and reduce the vanishing gradient problem. Over the past years, numerous unbounded activation functions were proposed for neural networks, with softplus Dugas et al. (2000) and rectified linear unit (ReLU) Glorot et al. (2011) activation function being the notable examples. These activation functions especially ReLU has pivotal role in improving the training and performance of deep learning models. ReLU has been a cornerstone of deep learning models due to its computational efficiency and effectiveness in addressing the vanishing gradient problem. Since then, several variants of ReLU were proposed including Leaky ReLU Maas et al. (2013), sigmoid linear unit Misra (2020) and exponential linear unit Clevert et al. (2016), each offering unique advantages for building deep learning-based applications.

### v5

@@ Abstract
Deep learning, a branch of artificial intelligence, is a data-driven method that uses multiple layers of interconnected units (neurons) to learn intricate patterns and representations directly from raw input data. Empowered by this learning capability, it has become a powerful tool for solving complex problems and is the core driver of many groundbreaking technologies and innovations. Building a deep learning model is challenging due to the algorithm’s complexity and the dynamic nature of real-world problems. Several studies have reviewed deep learning concepts and applications. However, the studies mostly focused on the types of deep learning models and convolutional neural network architectures, offering limited coverage of the state-of-the-art deep learning models and their applications in solving complex problems across different domains. Therefore, motivated by the limitations, this study aims to comprehensively review the state-of-the-art deep learning models in computer vision, natural language processing, time series analysis and pervasive computing. We highlight the key features of the models and their effectiveness in solving the problems within each domain. Furthermore, this study presents the fundamentals of deep learning, various deep learning model types and prominent convolutional neural network architectures. Finally, challenges and future directions in deep learning research are discussed to offer a broader perspective for future researchers.
@@
Deep Learning , Fundamentals , State-of-the-art Applications , Challenges
journal: Neurocomputing
@@
[1]organization=School of Computer Sciences, Universiti Sains Malaysia,addressline=Jalan Universiti, city=Gelugor, postcode=11800, state=Pulau Pinang, country=Malaysia
[…]
[2]organization=Department of Computer Science, Adekunle Ajasin University,addressline=Ikare Akoko-Ipeme Rd, city=Akungba-Akoko, postcode=P.M.B 001, state=Ondo State, country=Nigeria
@@ 1 Introduction
Deep learning has revolutionized many applications across a variety of industries and research. The application of deep learning can be found in healthcare [1], smart manufacturing [2], robotics [3], cybersecurity [4] etc., solving challenging and complex problems such as disease diagnosis, anomaly detection, object detection and malware attack detection. Deep learning is a subset of machine learning that focuses on learning from data using artificial neural networks with many layers, known as deep neural networks. An artificial neural network is a computational model that imitates the working principles of a human brain. The computational models are composed of an input layer which receives the input data, multiple processing layers that learn the representation of data and the output layer which produces the output of the model.
@@
Several studies have been conducted to discuss the concept and application of deep learning in the last few years, as listed in Table 1. The studies addressed or focused on several aspects of deep learning, such as types of deep learning models, learning approaches and strategies, convolutional neural network (CNN) architectures, deep learning applications and challenges. In [5], the authors provided fundamentals of deep learning and highlighted different types of deep learning models, such as convolutional neural networks, autoencoder and generative adversarial networks. Then, the applications of deep learning in various domains are discussed, and some challenges associated with deep learning applications are presented. Another survey [6] provided a comprehensive analysis of supervised, unsupervised and reinforcement learning approaches and compared the different learning strategies such as online, federated and transfer learning. Finally, the current challenges of deep learning and future direction are discussed.
In [7], the authors provided a comprehensive review of the popular CNN architectures used in computer vision tasks, highlighting their key features and advantages. Then, the applications of deep learning in medical imaging and the challenges are discussed. A similar survey is reported in [8], where the different supervised and unsupervised deep learning models are highlighted, and the popular CNN architectures are compared and discussed. In another survey [9], the authors focused on the applications of deep learning in computer vision, natural language processing and speech and audio processing. The different types of deep learning models are also discussed. In [10], the authors focused on the different types of deep learning models and provided a summary of deep learning applications in various domains.
@@
[5] A short review of the fundamentals of DL networks and discusses different types of neural networks, DL applications and challenges. Lack of analysis of CNN architectures and limited coverage of deep learning fundamentals.
[6] Discusses the learning approaches (supervised, unsupervised and reinforcement learnings), learning strategies, and DL challenges Lack of fundamentals of deep learning, CNN architectures and DL applications.
[7] Discusses different types of DL networks, CNN fundamentals and architectures, DL challenges and medical imaging applications Limited discussion on DL applications such as natural language processing and time series analysis.
[8] A short review of the fundamentals of neural networks and discusses different types of DL networks, CNN architectures and applications. Limited discussion on DL applications and no discussion of DL challenges.
[9] Discusses different types of DL networks and DL applications and challenges Lack of analysis of CNN architectures and limited coverage of deep learning fundamentals.
[10] Discusses different types of DL networks and provides a summary of DL applications Lack of fundamentals of deep learning, analysis of CNN architectures and limited discussion on DL applications.
@@ 2.1 Layers
A deep learning model is characterized by having numerous hidden layers. The hidden layers are responsible for learning and extracting complex features from the input data. A hidden layer is composed of an arbitrary number of neurons which serves as the fundamental building block of a neural network as shown in Fig. 1. A neuron consists of an arbitrary number of inputs, each associated with a weight, which controls the flow of information into the neuron during the forward pass. The flow of information, or forward pass, involves the computation of summation of the weighted input, followed by the application of a transformation function to the weighted sum. Let $\mathbf{x}={x_{1},x_{2},...,x_{d}},j=0,1,...,d$ be the input vector with $d$ dimensions and $\mathbf{w}_{i}^{l}$ denotes the weights that are connected to neuron $i$ in layer $l$. The forward pass to neuron $i$ in layer $l$ is defined as
$z_{i}^{l}=\mathbf{w}^{l}_{i}\cdot\mathbf{x}=\sum_{j=0}^{d}w_{i,j}^{l}\cdot x_{j}$ (1)
$a_{i}^{l}=g(z_{i}^{l})$ (2)
where $x_{j}$ is the input vector of size $d$, $w_{i,j}^{l}$ is the weight associated with $x_{j}$, connecting the input to neuron $i$ at layer $l$, and $g$ is the transformation function also known as activation function. It is worth noting that $w_{0}$ is called bias and $x_{0}=1$. A hidden layer wherein each neuron is connected to all neurons of the previous layer is known as fully-connected layer. The forward pass computation of a neuron $a_{i}^{l}$ in layer $l$ receiving a set of input from layer $l-1$ can be generalized as follows:
$a_{i}^{l}=g(\sum_{j=0}^{d}w_{i,j}^{l}\cdot a_{j}^{l-1})=\mathbf{w}^{l}_{i}\cdot\mathbf{a}^{l-1}$ (3)
where $\mathbf{a}^{l-1}$ is the input vector from layer $l-1$.
@@
Another crucial layer type is the convolutional layer, which is primarily used for processing data that has underlying structures, such as spatial patterns in image data or temporal patterns in time series data. Unlike fully-connected layer, each neuron in a convolutional layer is connected only to a subset of neurons in the previous layer as shown in Fig. 2. As shown in the figure, each neuron in the hidden layer is connected only to a local region, a subset of nine input neurons, and the weights are shared across the input data. The weight-sharing not only significantly reduces the number of parameters of the neural network, but also allows the network to learn the same features across different spatial locations in the input [11]. Let $\mathbf{x}$ be a two-dimensional input data with $H\times W\times C$ such as a grayscale image, where $H$ is the height, $W$ is the width and $C$ is the channel of the image data. The computation of a neuron in convolutional layer $l$ is defined as
$z_{i,j,d}^{l}=\sum_{m=0}^{k_{1}}\sum_{n=0}^{k_{2}}w_{m,n,d}^{l}\cdot x_{i\cdot s+m,j\cdot s+n,d}$ (4)
$a_{i,j,d}^{l}=g(z_{i,j,d}^{l})$ (5)
where $w_{m,n,d}$ is the weight connecting the input data $x_{i,j,d}$ within the window to the neuron in layer $l$ and $s$ is the stride or the step size of the window as it moves across the input data. This convolution operation where the filter is represented by $w_{m,n,d}$ slides over the input data $x_{i,j,d}$, producing a set of output values called feature map.
@@
Pooling layers are commonly applied after successive convolutional layers to progressively reduce the spatial dimensions of the feature maps. The spatial reduction is performed by computing the summary of a subset of the feature values in the feature maps as shown in Fig. 3. The pooling operation, as shown in the figure, can use either the maximum or average method with a pooling size of $2\times 2$ applied across the entire feature map, thereby reducing the size of the feature map. In addition to spatial reduction, pooling layers decrease the number of parameters and provide translation-invariant features [11]. Let $\mathbf{a}^{l}$ denotes feature map (hidden layer) $l$, a pooling operation with a pooling size of $m\times n$ is defined as
$\mathbf{a}^{l+1}=\texttt{pool}(\mathbf{a}^{l}_{i\cdot s+m,j\cdot s+n})$ (6)
where pool is either maximum or average function and $s$ is the stride or step size of the window as it moves across the feature map.
Figure 3: Summaries of the feature map using maximum or average pooling to produce a reduced feature map.
@@ 2.2 Attention Mechanisms
where $g$ is a composite function that performs a sequence of operations to generate the attention or the weights and $f$ applies the generated attention $g(x)$ on the input $x$.
@@
For instance, the squeeze-and-excitation (SE) attention generates the attention through five consecutive operations [12]. First, the input is vectorized using global average pooling. Then the vector is passed to two fully-connected layers, where the first one with ReLU activation and the second one with sigmoid activation. SE attention was a pioneer in channel attention. The attention module assigns varying weights to the channels of the feature maps. SE attention suffers from computational cost and the use of global average which may cause information loss at the spatial level. Several efforts have been made to improve SE attention. Global Second-order Pooling (GSoP) attention performs $1\times 1$ convolution on the feature maps to reduce the number of channels, and then computes the pairwise channel correlation which is used to generate the weights [13]. Efficient Channel Attention (ECA) replaced the fully-connected layers with 1d convolution to reduce the number of parameters and the computational cost [14].
@@
Temporal attention is an attention module that focuses on specific time steps in a sequence of data such time series and video (sequence of images). In video processing such as recognizing human actions, temporal attention is used to focus on key frames at different point in time that contains crucial information for predicting the ongoing activity. Temporal adaptive module (TAM) is a temporal attention that can focus on short-term (local) information and global context information of the data [15]. The composite function consists of local branch for generating attention weights and global branch for generating channel-wise adaptive kernel. First, the input feature map is squeezed using global average pooling to reduce the computational cost. Subsequently, the local branch executes two 1D convolution operations, with the first convolution using ReLU, and sigmoid activation for the second to generate the local weights. The local weights are then multiplied with the featuremap. Meanwhile, the global branch is composed of two fully-connected layers, with the first layer using ReLU and second layer employing softmax function to generate the adaptive kernel (weights). Self-attention is a form of temporal attention, initially proposed for machine translation to enable the deep learning models to attend different words in a sequence relative to other words [16]. The attention module has become a fundamental building block in various natural language processing applications. To generate the attention weights, the input (word embeddings) is transformed by linear projection to compute query, key and value. Then, dot product between query and key is computed, and the resultant is normalized by the square root of the size of the key. Finally, the attention weights are obtained by applying the softmax function. Self-attention is the fundamental building block of the Transformer architecture, a key deep learning model in natural language processing.
@@
Spatial attention focuses on specific regions or spatial location of the input data, enabling the deep learning models to selectively emphasize and ignore certain features. In the context of computer vision, spatial attention is crucial in capturing the spatial relationships and context within an image for accurate prediction. Attention gate is a spatial attention that can identify and focus the salient regions and suppress feature responses of the insignificant ones. The composite function consists of ReLU activation followed by $1\times 1$ convolution to reduce channel dimension of the feature maps to a singular feature map. Finally, sigmoid is applied to the feature map to generate attention weights [17]. The self-attention in the standard Transformer is not effective in handling image data due to its inherent sequential processing nature and lacks the ability to capture spatial dependencies and local patterns. To address this limitation, the Vision Transformer (ViT) treats images as a sequence of non-overlapping patches. A similar computational pipeline is used to generate the attention weights, the sequence of patches is transformed by linear projection to compute the query, key and value [18]. The same operations are employed to generate the attention weights. Self-attention is computationally costly due to its quadratic complexity especially dealing with image data. To reduce the complexity, two learnable linear layers, independent of the input data are adopted as the key and value vectors [19].
@@ 2.3 Activation Functions
The role of the activation function is to transform the weighted sum into a more classifiable form. This is crucial to the learning behaviour of the deep learning model, generating non-linear relationships between the input and the output of the model. The activation function, combined with many hidden layers allow the neural network to approximate highly complex, non-linear functions. Many activation functions are available for use in neural networks, and some of the functions are shown in Fig. 4 - Fig 6. The figures show the plot of the three popular activation functions. The sigmoid is a classic example of activation function which is used in logistic regression. Sigmoid activation function maps the weighted sum to a value in the range of 0 and 1 which can be used for classification. Hyperbolic tangent is another popular choice of bounded activation function which produces an output between -1 and 1. Hyperbolic tangent has a stronger gradient, hence the neural network training often converges faster than sigmoid [20]. For many years, sigmoid and hyperbolic tangent are the commonly used activation functions. Nevertheless, the activation functions suffer from vanishing gradient problem, hindering the efficient training of deep neural networks with many layers [21].
It was shown that neural networks with unbounded activation functions have the universal approximation property and reduce the vanishing gradient problem. Over the past years, numerous unbounded activation functions were proposed for neural networks, with softplus [22] and rectified linear unit (ReLU) [23] activation function being the notable examples. These activation functions especially ReLU has pivotal role in improving the training and performance of deep learning models. ReLU has been a cornerstone of deep learning models due to its computational efficiency and effectiveness in addressing the vanishing gradient problem. Since then, several variants of ReLU were proposed including Leaky ReLU [24], sigmoid linear unit [25] and exponential linear unit [26], each offering unique advantages for building deep learning-based applications.